An electron initially at rest passes from one plate of a charged capacitor to another. The voltage difference between plates is 10^7 V. Use (relativistic) conservation of energy to find the final velocity of the electron.
Answer #1

using the conservation of energy

kinetic energy = potential energy

0.5mv^2 = qV

0.5*9.11*10^-31*v^2 = 1.6*10^-19*10^7

v = 1.874*10^9 m/s

so the final velocity of the electron is 1.874*10^9 m/s